OpenStudy (anonymous):

if tanө= -5/12 and find cos ө (i need 2 see the working don`t know how to work this out but i have the answer)

OpenStudy (agreene):

\[\theta = \tan^{-1} (-5/12)\] \[\cos \theta = \cos ( \tan^{-1} (-5/12)) = 12/13\]